OpenStudy (superhelp101):

According to the periodic table, carbon has an atomic mass of 12.011 u. This indicates that the most abundant isotope of carbon is carbon-11. carbon-12. carbon-13. carbon-14.

OpenStudy (superhelp101):

@sweetburger any idea how to do this?

OpenStudy (superhelp101):

@GAzela1997

OpenStudy (superhelp101):

@DW101796

OpenStudy (superhelp101):

@lowcard2 @Luigi0210 @SarahEZZMcK @Abhisar @Awesome781 @anonymous_user @@dan815

OpenStudy (abhisar):

Which do you think it is? Why or why not? It would be easier to tutor you if you explained what about the question is giving you trouble.

OpenStudy (abhisar):

The most abundant will be that one whose atomic mass is closest to the average atomic mass

OpenStudy (abhisar):

can u select the correct option now ?

OpenStudy (anonymous):

I got carbon-12.

OpenStudy (superhelp101):

yeah i really did think it was carbon-12 too

OpenStudy (anonymous):

Do you understand how?

OpenStudy (superhelp101):

is it right?

OpenStudy (anonymous):

Yes.

OpenStudy (superhelp101):

not really can u explain please?

OpenStudy (superhelp101):

wait no i got it!!

OpenStudy (anonymous):

The most abundant will be that one whose atomic mass is closest to the average atomic mass

OpenStudy (anonymous):

Okay!

OpenStudy (superhelp101):

thanks! can u help me with other one please?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

I can try.

OpenStudy (superhelp101):

Which of the following atoms has the same number of neutrons as 19F? 16O 18F 20Ne 23Na how do u do this one @DW101796

OpenStudy (anonymous):

19 F has 19 - 9 = 10 neutrons 20 Ne = 20-10 = 10

OpenStudy (anonymous):

See how they both equal 10?

OpenStudy (superhelp101):

yes but why are you subtracting 19 by 9? is that the rule?

OpenStudy (anonymous):

Yes, for F.

OpenStudy (superhelp101):

oh so the answer would be the second one right?

OpenStudy (anonymous):

20Ne would be your answer. 3rd one.

OpenStudy (superhelp101):

ooooo yea i see now because that way both sides would equal 10. okay. Thanks!

OpenStudy (anonymous):

No problem man!